\begin{thebibliography}{23}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Abadi et~al.(2016)Abadi, Barham, Chen, Chen, Davis, Dean, Devin, Ghemawat, Irving, Isard, et~al.]{abadi2016tensorflow}
Mart{\'\i}n Abadi, Paul Barham, Jianmin Chen, Zhifeng Chen, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Geoffrey Irving, Michael Isard, et~al.
\newblock Tensorflow: a system for large-scale machine learning.
\newblock In \emph{Osdi}, volume~16, pp.\  265--283. Savannah, GA, USA, 2016.

\bibitem[Bradbury et~al.(2018)]{jax2018github}
James Bradbury et~al.
\newblock {JAX}: composable transformations of {P}ython+{N}um{P}y programs, 2018.
\newblock URL \url{http://github.com/google/jax}.

\bibitem[Cai et~al.(2021)Cai, Mao, Wang, Yin, and Karniadakis]{cai2021physics}
Shengze Cai, Zhiping Mao, Zhicheng Wang, Minglang Yin, and George~Em Karniadakis.
\newblock Physics-informed neural networks (pinns) for fluid mechanics: A review.
\newblock \emph{Acta Mechanica Sinica}, 37\penalty0 (12):\penalty0 1727--1738, 2021.

\bibitem[Chiu et~al.(2022)Chiu, Wong, Ooi, Dao, and Ong]{chiu2022can}
Pao-Hsiung Chiu, Jian~Cheng Wong, Chinchun Ooi, My~Ha Dao, and Yew-Soon Ong.
\newblock Can-pinn: A fast physics-informed neural network based on coupled-automatic--numerical differentiation method.
\newblock \emph{Computer Methods in Applied Mechanics and Engineering}, 395:\penalty0 114909, 2022.

\bibitem[Cho et~al.(2023)Cho, Nam, Yang, Yun, Hong, and Park]{cho2023separable}
Junwoo Cho, Seungtae Nam, Hyunmo Yang, Seok-Bae Yun, Youngjoon Hong, and Eunbyung Park.
\newblock Separable physics-informed neural networks.
\newblock \emph{Advances in Neural Information Processing Systems}, 2023.

\bibitem[Gao et~al.(2023)Gao, Köhler, and Foster]{folx}
Nichoals Gao, Jonas Köhler, and Adam Foster.
\newblock folx - forward laplacian for jax, 2023.
\newblock URL \url{http://github.com/microsoft/folx}.

\bibitem[Han et~al.(2018)Han, Jentzen, and E]{han2018solving}
Jiequn Han, Arnulf Jentzen, and Weinan E.
\newblock Solving high-dimensional partial differential equations using deep learning.
\newblock \emph{Proceedings of the National Academy of Sciences}, 115\penalty0 (34):\penalty0 8505--8510, 2018.

\bibitem[Han et~al.(2019)Han, Zhang, and E]{han_solving_2019}
Jiequn Han, Linfeng Zhang, and Weinan E.
\newblock Solving many-electron schr{\"o}dinger equation using deep neural networks.
\newblock \emph{Journal of Computational Physics}, 399:\penalty0 108929, December 2019.
\newblock ISSN 00219991.
\newblock \doi{10.1016/j.jcp.2019.108929}.
\newblock URL \url{http://arxiv.org/abs/1807.07014}.

\bibitem[He et~al.(2023)He, Li, Shi, Gao, Zhang, Bian, Wang, and Liu]{he2023learning}
Di~He, Shanda Li, Wenlei Shi, Xiaotian Gao, Jia Zhang, Jiang Bian, Liwei Wang, and Tie-Yan Liu.
\newblock Learning physics-informed neural networks without stacked back-propagation.
\newblock In \emph{International Conference on Artificial Intelligence and Statistics}, pp.\  3034--3047. PMLR, 2023.

\bibitem[Hermann et~al.(2020)Hermann, Sch{\"a}tzle, and No{\'e}]{hermann2020deep}
Jan Hermann, Zeno Sch{\"a}tzle, and Frank No{\'e}.
\newblock Deep-neural-network solution of the electronic {Schr{\"o}dinger} equation.
\newblock \emph{Nature Chemistry}, 12\penalty0 (10):\penalty0 891--897, 2020.

\bibitem[Hu et~al.(2023)Hu, Shi, Karniadakis, and Kawaguchi]{hu2023hutchinson}
Zheyuan Hu, Zekun Shi, George~Em Karniadakis, and Kenji Kawaguchi.
\newblock Hutchinson trace estimation for high-dimensional and high-order physics-informed neural networks.
\newblock \emph{arXiv preprint arXiv:2312.14499}, 2023.

\bibitem[Kovachki et~al.(2021)Kovachki, Li, Liu, Azizzadenesheli, Bhattacharya, Stuart, and Anandkumar]{kovachki2021neural}
Nikola Kovachki, Zongyi Li, Burigede Liu, Kamyar Azizzadenesheli, Kaushik Bhattacharya, Andrew Stuart, and Anima Anandkumar.
\newblock Neural operator: Learning maps between function spaces.
\newblock \emph{arXiv preprint arXiv:2108.08481}, 2021.

\bibitem[LeCun et~al.(2015)LeCun, Bengio, and Hinton]{lecun2015deep}
Yann LeCun, Yoshua Bengio, and Geoffrey Hinton.
\newblock Deep learning.
\newblock \emph{nature}, 521\penalty0 (7553):\penalty0 436--444, 2015.

\bibitem[Li et~al.()Li, Ye, Jiang, Wen, Wang, Li, Li, He, Chen, Ren, et~al.]{li2023forward}
Ruichen Li, Haotian Ye, Du~Jiang, Xuelan Wen, Chuwei Wang, Zhe Li, Xiang Li, Di~He, Ji~Chen, Weiluo Ren, et~al.
\newblock A computational framework for neural network-based variational monte carlo with forward laplacian.
\newblock \emph{Nature Machine Intelligence}.
\newblock URL \url{https://www.nature.com/articles/s42256-024-00794-x}.

\bibitem[Meng et~al.(2021)Meng, Song, Li, and Ermon]{meng2021estimating}
Chenlin Meng, Yang Song, Wenzhe Li, and Stefano Ermon.
\newblock Estimating high order gradients of the data distribution by denoising.
\newblock \emph{Advances in Neural Information Processing Systems}, 34:\penalty0 25359--25369, 2021.

\bibitem[Pang et~al.(2020)Pang, Xu, Li, Song, Ermon, and Zhu]{pang2020efficient}
Tianyu Pang, Kun Xu, Chongxuan Li, Yang Song, Stefano Ermon, and Jun Zhu.
\newblock Efficient learning of generative models via finite-difference score matching.
\newblock \emph{Advances in Neural Information Processing Systems}, 33:\penalty0 19175--19188, 2020.

\bibitem[Paszke et~al.(2017)Paszke, Gross, Chintala, Chanan, Yang, DeVito, Lin, Desmaison, Antiga, and Lerer]{paszke2017automatic}
Adam Paszke, Sam Gross, Soumith Chintala, Gregory Chanan, Edward Yang, Zachary DeVito, Zeming Lin, Alban Desmaison, Luca Antiga, and Adam Lerer.
\newblock Automatic differentiation in pytorch.
\newblock 2017.

\bibitem[Pfau et~al.(2020)Pfau, Spencer, de~G.~Matthews, and Foulkes]{ferminet}
D.~Pfau, J.S. Spencer, A.G. de~G.~Matthews, and W.M.C. Foulkes.
\newblock Ab-initio solution of the many-electron schr{\"o}dinger equation with deep neural networks.
\newblock \emph{Physical Review Research}, 2:\penalty0 033429, 2020.
\newblock \doi{10.1103/PhysRevResearch.2.033429}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevResearch.2.033429}.

\bibitem[Raissi et~al.(2019)Raissi, Perdikaris, and Karniadakis]{raissi2019physics}
Maziar Raissi, Paris Perdikaris, and George~E Karniadakis.
\newblock Physics-informed neural networks: A deep learning framework for solving forward and inverse problems involving nonlinear partial differential equations.
\newblock \emph{Journal of Computational Physics}, 378:\penalty0 686--707, 2019.

\bibitem[Rogallo \& Moin(1984)Rogallo and Moin]{rogallo1984numerical}
Robert~S Rogallo and Parviz Moin.
\newblock Numerical simulation of turbulent flows.
\newblock \emph{Annual review of fluid mechanics}, 16\penalty0 (1):\penalty0 99--137, 1984.

\bibitem[Sirignano \& Spiliopoulos(2018)Sirignano and Spiliopoulos]{sirignano2018dgm}
Justin Sirignano and Konstantinos Spiliopoulos.
\newblock Dgm: A deep learning algorithm for solving partial differential equations.
\newblock \emph{Journal of computational physics}, 375:\penalty0 1339--1364, 2018.

\bibitem[Wang et~al.(2022)Wang, Li, He, and Wang]{wang20222}
Chuwei Wang, Shanda Li, Di~He, and Liwei Wang.
\newblock Is ${L}^2$ physics informed loss always suitable for training physics informed neural network?
\newblock In \emph{Advances in Neural Information Processing Systems}, 2022.
\newblock URL \url{https://openreview.net/forum?id=cy1TKLRAEML}.

\bibitem[Ye et~al.(2023)Ye, Li, and Jiang]{LapJAX}
Haotian Ye, Ruichen Li, and Du~Jiang.
\newblock Lapjax, 2023.
\newblock URL \url{https://github.com/YWolfeee/lapjax}.

\end{thebibliography}

\documentclass{article} 
\usepackage{arxiv,times}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\def\rvg{{\mathbf{g}}}

\def\rvu{{\mathbf{i}}}

\def\rvu{{\mathbf{u}}}

\def\rvx{{\mathbf{x}}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{marvosym}
\usepackage{mathrsfs}
\usepackage{blindtext}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{example}[theorem]{Example}

\usepackage{hyperref}
\hypersetup{hidelinks}
\usepackage{url}
\usepackage{enumerate}
\usepackage{cleveref}
\usepackage{booktabs}
\usepackage{tabularx}
\usepackage{multirow}
\usepackage{algorithm}
\usepackage{algpseudocode}
\newcommand{\pa}[1]{\partial_{#1}}
\newcommand{\pab}[2]{\frac{\partial #1}{\partial #2}}
\newcommand{\ppabc}[3]{\frac{\partial^2 #1}{\partial #2\partial #3}}
\newcommand{\bgeq}[1]{\begin{equation}#1\end{equation}}
\newcommand{\doph}{DOF\xspace}

\newcommand{\fl}{Forward Laplacian\xspace}

\newcommand{\name}{LapNet\xspace}

\newcommand{\lrc}[1]{{\textcolor{blue}{[lrc: #1]}}}
\newcommand{\yht}[1]{{\textcolor{purple}{[HT: #1]}}}

\title{\doph: Accelerating High-order Differential Operators with Forward Propagation}

\author{%
  Ruichen Li$^{1,}$\thanks{Equal contributions.} \quad Chuwei Wang$^{2,*}$ \quad Haotian Ye$^{3,*}$  \quad Di He$^{1,}$\thanks{Corresponding to: \texttt{\{dihe,wanglw\}@pku.edu.cn}.} \quad Liwei Wang$^{1,4,\dag}$\\
  $^1$National Key Laboratory of General Artificial Intelligence, Peking University\\
  $^2$California Institute of Technology\quad
  $^3$Stanford University\\
  $^4$Center for Machine Learning Research, Peking University\\
}

\newcommand{\hes}{\textit{Hessian-based} methods }

\iclrfinalcopy 
\begin{document}

\maketitle

\begin{abstract}
Solving partial differential equations (PDEs) efficiently is essential for analyzing complex physical systems. Recent advancements in leveraging deep learning for solving PDE have shown significant promise. However, machine learning methods, such as Physics-Informed Neural Networks (PINN), face challenges in handling high-order derivatives of neural network-parameterized functions. Inspired by Forward Laplacian, a recent method on accelerating Laplacian computation, we propose an efficient computational framework, Differential Operator with Forward-propagation (DOF), for calculating general second-order differential operators without losing any precision. We provide rigorous proof of the advantages of our method over existing methods, demonstrating two times improvement in efficiency and reduced memory consumption on \textit{any} architectures. Empirical results illustrate that our method surpasses traditional automatic differentiation (AutoDiff) techniques, achieving 2x improvement on the MLP structure and nearly 20x improvement on the MLP with Jacobian sparsity. 
\end{abstract}

\section{Introduction}

Partial differential equations (PDEs) play a pivotal role in understanding and predicting the behavior of physical systems. While classical numerical methods have proven effective in some cases, they can be prohibitively challenging when dealing with complicated problems, e.g., turbulence in fluid dynamics \citep{rogallo1984numerical} and high dimensional equations \cite{han2018solving}. Recently, the advent of deep learning\cite{lecun2015deep} has spurred a wave of innovations in leveraging neural networks (NN) for numerical solutions of PDEs. These NN-based approaches have been applied to various problems, including fluid dynamics \cite{cai2021physics,kovachki2021neural}, high-dimensional optimal control problems \cite{wang20222,han_solving_2019}, and quantum many-body problems\cite{ferminet,hermann2020deep}. Notable works, such as the Physics-Informed Neural Network (PINN)\cite{raissi2019physics}, showcase the potential of neural networks in capturing the underlying physics of systems governed by PDEs. 

Among these works, the central idea is to parameterize the solution as a neural network, and optimize this expressive network with the guidance from PDE. The ever-growing \textit{AutoDiff} packages\cite{jax2018github,paszke2017automatic,abadi2016tensorflow}  enables convenient calculation of associated quantities such as residual losses and derivatives, avoiding discretization errors in classical methods. However, unlike most computer vision or natural language processing tasks in which first-order derivatives are sufficient for optimizations, in PDE-relevant problems one has to deal with high-order derivatives.
This raises a significant challenge as common AutoDiff packages are computationally intensive under this circumstance\cite{pang2020efficient,meng2021estimating}.
There have been attempts to address this issue \cite{he2023learning,hu2023hutchinson,sirignano2018dgm,chiu2022can}. However, they resort to either randomized methods\cite{he2023learning,hu2023hutchinson} or numerical differentiation\cite{sirignano2018dgm,chiu2022can}, which introduce unsatisfying statistical errors and are limited in problems where high precision is not demanded. 

Recently, \citet{li2023forward} proposed a computational framework, Forward Laplacian (FL). This framework is designed specifically for accelerating Laplacian operator computation and thus can significantly boost the computation of Laplacian-relevant PDE like Schrödinger equation in quantum chemistry\cite{ferminet}. Remarkably, Forward Laplacian is proven to be precision-preserved as it introduces no statistical errors at all. 
Consequently, it is natural to ask whether the idea of FL can be leveraged for other PDEs associated with high-order derivatives computation.

To fill this gap, we develop a new computational framework named Differential Operator with Forward-propagation (\doph). \doph shares a similar computational procedure as FL, yet can be applied to compute general second-order differential operators. We demonstrate that \doph outperforms conventional AutoDiff methods in memory and computational cost by a large margin, both theoretically and practically. In practice, \doph can accelerate NN-based solvers across a wide range of PDEs such as the non-homogeneous heat equation and Klein-Gordon equation.

The contribution of this paper is summarized as follows:
\begin{itemize}
\item We generalize the Forward Laplacian method and propose \doph to \textit{precisely} compute arbitrary second-order differential operators of neural networks (\doph).
\item We demonstrate that \doph improves computation efficiency and memory consumption simultaneously, regardless of the architecture of neural networks, both theoretically and empirically. The improvement can be significant in common architectures like MLP.
\end{itemize}

\section{Method}
In PINNs and many other NN-based PDE solvers, the solution of a PDE, $\phi(\rvx)$, is parameterized as neural networks, $\phi(\rvx):=\phi(\rvx;\theta)$, where $\theta$ represent the NN parameters. These methods necessitate computing the high-order derivatives of a neural network. It has been shown that standard \textit{AutoDiff} methods are not efficient for the high-order derivatives calculation\cite{pang2020efficient,meng2021estimating}.

In this work, we focus on the calculation of the
second-order differential operators, which have the form 
\bgeq{\label{general_L}
\mathcal{L}: \phi(\rvx)\to \sum_{1\leq i,j\leq N}a_{ij}(\rvx)\partial^2_{ij}\phi(\rvx)+\sum_{i\leq N} b_i(\rvx)\partial_i \phi(\rvx)+c(\rvx)\phi(\rvx).
}
Here, $a_{ij}, b_i, c: \mathbb{R}^N\rightarrow \mathbb{R}$ are coefficients in the second order operator, and $N$ is the input dimension (the time variable is comprised in $\rvx$ for evolution equations). In practice, the first term dominates the computation cost. Thus, for brevity and clarity, we will only focus on the case when $b_i\equiv c\equiv 0$ in the following discussions. We always denote the symmetric matrix $\big(a_{ij}(\rvx)\big)_{i,j}$ as $A(\rvx)$.

\subsection{Forward Laplacian}
In standard AutoDiff packages, the second-order operator calculation is based on the Hessian matrix $H=(\pa{ij}\phi(x))_{i,j}$. We call those methods \textit{Hessian-based} methods. They use multiple Jacobian calculations to derive the Hessian matrix, resulting in a huge computation cost. Recently, \citet{li2023forward} proposed a new computational framework, Forward Laplacian (FL), which primarily focuses on accelerating the calculation of Laplacian, i.e., $A\equiv I_{N}$. Below we briefly review this method.

FL computes the Laplacian with \textit{one} efficient forward pass, avoiding redundant calculation in the Hessian-based approach. Following the notation in \citet{li2023forward}, we describe FL in a computation graph $\mathcal{G}$. The node set $V=\{v^i|i=0,1,...,M\}$ represents the operations or variables used in a neural network. We use the abbreviation $i\to j$ if there is a directed edge from $v^i$ to $v^j$ in $\mathcal{G}$, and we denote operations as $F$, e.g., $v^j=F_j(\{v^i:i\to j\})$ for all $j\geq 0$. Notice that the node indices are arranged according to the topological order, i.e., for all $i\to j$, we have $i<j$. The output of $\phi$ is denoted by $v^M$. In addition, $\nabla$ and $\Delta$ represent the gradient operator and Laplacian operator with respect to the input, respectively. Detailed notations can be found in the \cref{apdxA}.

Specifically, according to node dependency, FL sequentially computes the Laplacian tuple $(v^i,\nabla v^i, \Delta v^i)$ associated with each node. In a simplified case where $v^i$ depends only on $v^{i-1}$, i.e., $v^i=F_i(v^{i-1})$, the graph can be represented as a chain. We can compute the output tuple in a forward pass:
\begin{equation}
    (\rvx,\nabla \rvx, \Delta \rvx) \rightarrow (v^0,\nabla v^0, \Delta v^0) \rightarrow \cdots \rightarrow (v^M,\nabla v^M, \Delta v^M)
\end{equation}

The propagation rule of Laplacian tuple is derived through the chain rule:
\begin{equation}
\label{eq:chain}
    v^{i}=F_{i}(v^{i-1}),~ \nabla v^{i}=\pa{v^{i-1}}F_{i}\nabla v^{i-1}, ~ \Delta v^{i}=\pa{v^{i-1}}^2F_{i}|\nabla v^{i-1}|^2 + \pa{v^{i-1}}F_{i}\Delta v^{i-1}
\end{equation}

For the general computation graph, we can generalize \cref{eq:chain} to the following formula:
\begin{align}
        v^j &= F_j(\{v^i:i\to j\})\label{eq:fl_method1} \\
        \nabla_{ } v^{j} &= \sum\limits_{i:i\to j}\pab{F_{j}}{v^i}\nabla_{ } v^{i} \label{eq:fl_method2}\\
        \Delta v^{j} &= \sum\limits_{\substack{i,l\\i\to j\ l\to j}}\ppabc{F_{j}}{v^{i}}{v^{l}}\nabla v^{i}\cdot \nabla v^{l} + \sum\limits_{i:i\to j} \pab{F_{j}}{v^{i}}\Delta v^{i}\label{eq:fl_method3}.
\end{align}
Then we can sequentially compute the Laplacian tuple for each node in the topological order according to \cref{eq:fl_method1,eq:fl_method2,eq:fl_method3}. As shown in \citet{li2023forward}, this approach outperforms the Hessian-based approach in numerous types of computational graphs. It has been successfully applied to solving the Schrodinger equation, resulting in over a magnitude of acceleration.


\subsection{\doph for the Second-order Differential Operators}

We now formally propose \doph to efficiently compute all kinds of second-order operators. For brevity, we denote $A(\rvx)$ as $A$ and $a_{ij}(\rvx)$ as $a_{ij}$ in the following discussion. To compute $\sum_{i,j}a_{ij}\pa{i}\pa{j}\phi(\rvx)$, we first decompose the coefficient matrix $A$ into $L^\top DL$ such that $D$ is a diagonal matrix whose diagonal elements are all $\pm 1$ and $0$. As $A$ is a symmetric matrix, this decomposition can be done easily. 
For instance, we can eigen-decompose $A=S^\top\Sigma S$, where $S$ is an orthogonal matrix and $\Sigma$ is the diagonal eigenvalue matrix, and choose $L=|\Sigma|^{1/2}S$ and $D=\text{sgn}(\Sigma)$. 

During the computation, for each node $v^k$, we compute tuple $(v^j,{\mathbf{g}}^j,s^j):=(v^j, L\nabla v^j, \mathcal{L}v^j)$. We can derive the following formula by chain rule:
\begin{align}
    \label{forward:func_elp10}
        v^j &= F_j(\{v^i:i\to j\}) \\
        {\mathbf{g}}^j &= \sum\limits_{i:i\to j}\pab{F_{j}}{v^i}{\mathbf{g}}^i \label{forward:func_elp20}\\
        s^{j} &= \sum\limits_{\substack{i,l\\i\to j\ l\to j}}\ppabc{F_{j}}{v^{i}}{v^{l}}{\mathbf{g}^{i}}^{\top}D \mathbf{g}^{l} + \sum\limits_{i:i\to j} \pab{F_{j}}{v^{i}}s^{i}\label{forward:func_elp30}
\end{align}
We can derive $\mathcal{L}\phi$ by sequentially applying this propagation rule to each node in the topological order. We will prove that this method outperform the Hessian-based methods in both the memory usage and computation cost for any neural network architecture: 

\begin{theorem}\label{thm1}
    The computation cost (counted in FLOPs) of \doph is at most half that of Hessian-based methods for any neural network architecture.
\end{theorem}
 \begin{theorem}\label{thm2}
    The memory consumption of \doph ($\mathcal{M}_1$) is smaller than that of Hessian-based methods ($\mathcal{M}_2$) for any neural network architecture.

    Specifically, $\mathcal{M}_1\lesssim\frac 2 L \mathcal{M}_2$ for an $L$-layer MLP.
\end{theorem}

The proof for \cref{thm1,thm2} could be found in \cref{abdx_b,apdx_c}. We remark that the memory and computation consumption of \doph can be further reduced in some specific network architectures. See \cref{sec:SMLP} for details.


To better understand the \doph method, we discuss the implementation of \doph on two special classes of second-order operators.

\paragraph{Elliptic Operator.} For elliptic operator, the coefficient matrix $A$ is positive so we have $D=I_{N}$. As a result, \cref{forward:func_elp10,forward:func_elp20,forward:func_elp30} are reduced to \cref{eq:fl_method1,eq:fl_method2,eq:fl_method3}. The only difference between \doph and Forward Laplacian here is the initial value of the tuple (i.e. the tuple at node $v^j$ for $j=0$). Thus, we utilize some existing Forward Laplacian package\cite{LapJAX,folx} to compute the elliptic operator.

\paragraph{Low-rank Coefficient Matrix.} It is well-known that computing the Hessian-vector product can be faster than computing the entire Hessian matrix in the standard AutoDiff package. Namely, computing Hessian-vector product takes $\mathcal{O}(1/N)$ cost of computing the entire Hessian matrix. Thus, in the standard AutoDiff package, if the coefficient matrix is a low-rank matrix, the calculation of the second-order operator can be accelerated through multiple Hessian-vector products. Similarly, the \doph method also accomplishes this acceleration when dealing with a low-rank coefficient matrix. 

If the coefficient matrix $A$ is a rank-$r$ matrix, we can eliminate the columns and rows associated with the zero eigenvalues in $L$ and $D$. Then we have $L'\in \mathbb{R}^{r\times N}$ and $D'\in \mathbb{R}^{r\times r}$ that still satisfy $A=L'^{\top}D'L'$. Thus, while the propagation rule of \doph remains the same, the dimension of ${\mathbf{g}}^k$ is reduced from $N$ to $r$. According to the analysis in the \cref{abdx_b,apdx_c}, this reduction in the dimension will lead to an $\mathcal{O}(r/N)$ reduction in both the memory usage and the computation cost.

\section{Results\label{sec:Results}}
\begin{table}[t]
\caption{Comparison between \doph and Hessian-based method on the MLP\label{tab:MLP}}
\begin{center}
\newcolumntype{Y}{>{\centering\arraybackslash}X}
\begin{tabularx}{\textwidth}{@{}lYYYYYY@{}}
\toprule\midrule
\multirow{2}{*}{Operator} & \multicolumn{3}{c}{GPU Memory Usage (MB)} & \multicolumn{3}{c}{Time (ms)} \\ 
                  &   Hessian    &   \doph    &   ratio    &   Hessian    &   \doph    &   ratio   \\ \midrule
                  Elliptic&   10421    &   3165    &   3.3    &   196.7    &  106.6    &    1.8  \\
                  Low-rank&    10427   &  2141     &    4.9   &    196.2   &    55.8    &   3.5  \\
                  General&   10429    &   3181    &   3.3    &   197.4    &   122.2    &  1.6    \\ \bottomrule
\end{tabularx}
\end{center}
\end{table}
\begin{table}[t]
\caption{Comparison between \doph and Hessian-based method on the MLP with Jacobian sparsity\label{tab:SMLP}}
\begin{center}
\newcolumntype{Y}{>{\centering\arraybackslash}X}
\begin{tabularx}{\textwidth}{@{}lYYYYYY@{}}
\toprule\midrule
\multirow{2}{*}{Operator} & \multicolumn{3}{c}{GPU Memory Usage (MB)} & \multicolumn{3}{c}{Time (ms)} \\ 
                  &   Hessian    &   \doph    &   ratio    &   Hessian    &   \doph    &   ratio   \\ \midrule
                  Elliptic&    21401   &   997    &   21.5    &   366.4    &   18.9    &   19.4   \\
                  Low-rank&   21401    &   869    &   24.6    &   366.2    &  12.7     &  28.9    \\
                  General&   21401    &   997    &    21.5   &     366.6    &  18.9  & 19.4 \\ \bottomrule
\end{tabularx}
\end{center}
\end{table}
In this section, we compare \doph with the standard Hessian-based approach on different operators and network architectures. We study three kinds of second-order operators: elliptic operator, elliptic operator with low-rank coefficient matrix, and general operator. As for the architecture, we choose the standard MLP structure and the MLP with Jacobian sparsity. Details could be found in \cref{apdx:exp_details}.

\subsection{Comparison on the MLP} The benchmark results for the MLP are shown in \cref{tab:MLP}. For both the elliptic and general operators, \doph demonstrates significant efficiency improvements, halving the computation cost and reducing memory usage by a third compared to the Hessian-based method, aligning with our theoretical predictions. In the case of the low-rank operator, \doph achieves even greater acceleration, attributed to the dimensionality reduction of $\mathbf{g}$. 
\subsection{Comparison on the MLP with Jacobian Sparsity\label{sec:SMLP}} As discussed in \citet{li2023forward}, the cost of Forward Laplacian method is significantly reduced when the Jacobian of intermediate component is sparse. We notice that this sparse Jacobian property also exists in some advanced PINN network architecture \cite{cho2023separable}. As a forward AutoDiff method, \doph can also leverage this property to significantly accelerate the calculation. The memory and time comparison are listed in \cref{tab:SMLP} and the architecture details can be found in \cref{apdx:exp_details}. Compared with the Hessian-based method, \doph can significantly reduce both the memory and computation consumption in the MLP with Jacobian sparsity, showing a great potential in applying our method to the advanced machine learning-based PDE solver.

\bibliography{iclr2024_workshop}
\bibliographystyle{iclr2024_workshop}
\newpage
\appendix
\section{Preliminary}\label{apdxA}

In this section, we briefly review computation framework in deep learning and auto differentiation, and introduce the notations we will use in the following sections.

\subsection{Computation Graph}
The computation graph serves as a descriptive language of deep learning models across various deep learning toolkits, including PyTorch \cite{paszke2017automatic}, TensorFlow \cite{abadi2016tensorflow}, and Jax \cite{jax2018github}. 

In the computation graph $\mathcal{G}$ associated with a function $\phi(\rvx),\ \rvx=(x_1,...x_N)\in\mathbb{R}^N$(we will always use $N$ for input dimension), the edges represent function arguments, and nodes represent operations or variables. Note that $\mathcal{G}$ is always a directed acyclic graph. We use $\{{v}^{-1},...{v}^{-N}\}$ to represent the external nodes of $\mathcal{G}$ (i.e., the input $\rvx$ of the neural network-parameterized function $\phi$), and use $\{v^0,... v^{M}\}$ to represent the internal nodes, sorted in topological orders. A node can be referred to as a neuron in the network. Specifically, we use $v^{M}$ to serve as the network output $\phi$. We use the abbreviation $i\to j$ if there is a directed edge from $v^i$ to $v^j$ in  $\mathcal{G}$. Furthermore, we denote operations as $F$, e.g., $v^j=F_j(\{v^i:i\to j\})$ for all $j\geq 0$.

\begin{example}\label{mlp_notatino}
Take Multi-Layer Perceptron (MLP) function $\phi(\rvx)$ as an example.

$\phi$ has the form $\phi=F_{L}\circ F_{L-1}\circ ...\circ F_0$, where $F_l=(F_{l,1},...F_{l,N_{l+1}}):\mathbb{R}^{N_{l}}\to\mathbb{R}^{N_{l+1}}$ are the mapping in each layer with $N_0=N$ and $N_{L+1}=1$. Let $\rvu^l=(u_1^l,u_2^l,...u^l_{N_l})\in\mathbb{R}^{N_l}$ be the vector consist of the neurons in the $l$-th layer, we have $\rvu^0=\rvx$ and $\rvu_{l+1}=F_l(\rvu^l)=\sigma(W^l\rvu^l+b^l)$ where $W^l\in\mathbb{R}^{N_{l+1}\times N_l},\ b^l\in\mathbb{R}^{N_{l+1}}$ are network parameters and $\sigma$ is a nonlinear function operated element-wise.

In this setting, the nodes $\{v^i\}_i$ in the computation graph are
\begin{align}
    x_1,...x_N,\ u^1_1,...u^1_{N_1},\  u^{1.5}_1,...u^{1.5}_{N_1},\ u^2_1,...,u^2_{N_2},...,u^{{L+1}}_1(=\phi(\rvx)),
\end{align}
and the operations generating the node representing $u_i^l$ and $u_i^{l-0.5}$ are $ u_i^l=\sigma(u_i^{l-0.5})$ and\\
$u^{l-0.5}_i=\sum_{j=1}^{N_{l-1}}W^{l-1}_{ij}u_j^{l-1} +b^{l-1}_i$, respectively.
\end{example}

\subsection{Auto Differentiation}
In most machine learning toolkits, auto differentiation(AutoDiff) implemented with back propagation algorithm is applied to compute the derivatives of neural network functions.

For a function $\phi(\rvx)$, this method first performs forward propagation to obtain the value of each variable $v^j$, and construct the computation graph $\mathcal{G}$.
Next, a backward process is employed by creating a new computation graph $\hat{\mathcal{G}}$ where node $\hat{v}^i\in \hat{\mathcal{G}}$ represents the operation to calculate $\pab{\phi}{v^i}$, $i=M,...,-N$. The associated computations are
\begin{equation}
\label{eq: backwarddd}
    \pab{\phi}{v^i} = \sum\limits_{j:i\to j} \pab{F_j}{v^i} \pab{\phi}{v^j},\ i=M-1,...,-N.
\end{equation}


\clearpage
\section{Proof of \cref{thm1}}\label{abdx_b}
Our proof follows similar idea to what is discussed in section 4.2 in \citet{li2023forward}, except that we give the proof for general second order operator here.

Recall that we are going to compute \cref{general_L} and we only need to analyze the computation cost of its first term. In the following, we will use the notation described in \cref{apdxA} for computation graphs.
The previous auto differentiation method obtains $\mathcal{L}\phi(\rvx)$ through computing the Hessian matrix and its inner product with $A$. It first performs forward propagation to obtain the value of each variable $v^j$. Next, a standard backward process is employed by creating a new computation graph $\hat{\mathcal{G}}$ where node $\hat{v}^i\in \hat{\mathcal{G}}$ represents the operation to calculate $\pab{\phi}{v^i}$, $i=M,...,-N$. The associated computations are
\begin{equation}
\label{eq: backward}
    \pab{\phi}{v^i} = \sum\limits_{j:i\to j} \pab{F_j}{v^i} \pab{\phi}{v^j},\ i=M-1,...,-N.
\end{equation}

The Hessian matrix is then obtained through the following forward-mode Jacobian calculation along $\mathcal{G}$ and $\hat{\mathcal{G}}$, respectively:
\begin{equation}
\label{eq: forward grad}
    \begin{aligned}
    \nabla_{ } v^{i} &= \sum\limits_{j:j\to i}\pab{F_{i}}{v^j}\nabla_{ } v^{j},\ i=-N,...M \quad 
    \end{aligned}
\end{equation}
\begin{equation}
\label{eq: hessian second}
    \begin{aligned}
    \nabla \hat v^i = \nabla_{ } \pab{\phi}{v^{i}}  =& \sum\limits_{\substack{j,l\\ i\to j\\ l\to j}} \ppabc{F_{j}}{v^{l}}{v^{i}}\pab{\phi}{v^{j}}\nabla_{ } v^{l} + \sum_{j:i\to j}\pab{F_{j}}{v^{i}}\nabla_{ }\pab{\phi}{v^{j}}, \ i=M-1,...-N\\
\end{aligned}
\end{equation}
where we always use $\nabla$ to denote $\nabla_\rvx$ for simplicity.

The bottleneck of Hessian computation comes from \cref{eq: forward grad} and \cref{eq: hessian second}. 

For every connected nodes $(j\to i)$, it takes $N$ float multiplication to yield $\pab{F_{i}}{v^j}\nabla_{ } v^{j}$. Therefore
\cref{eq: forward grad} takes $N|E|$ floating point operations (FLOPs) if we only count multiplications, where $E$ denotes the set of edges in $\mathcal{G}$. For \cref{eq: hessian second}, the second term also takes $N|E|$ FLOPs. To calculate the computational cost of the first term in \cref{eq: hessian second}, we first introduce two notations, $T$ and $R$, which are both sets of ordered tuples:
\begin{equation}
    \begin{aligned}
        T &= \{(i,l,j)| i\to j ,\ l \to j,\ \ppabc{F_j}{v^i}{v^l}\neq 0\},\\
        R &= \{(i,l)| \exists j\ s.t.\ (i, l, j)\in T\}.
    \end{aligned}
\end{equation}
The AutoDiff method sums over $j$ first to obtain $\sum\limits_{j:i\to j,l\to j} \ppabc{F_{j}}{v^{l}}{v^{i}}\pab{\phi}{v^{j}}$ for all $(i,l)\in R$, and then sums over $l$. 
For the first step, by leveraging the symmetry of Hessian matrix, it spends 1 FLOPS for a pair of $(i,l,j)$ and $(l,i,j)$ in $T$. Thus it spends $0.5|T|$ FLOPs in total. For the second step, for any $(i,l)\in R$, it takes $N$ FLOPs to multiply a vector $\nabla v^l$ with a scalar. Thus this steps takes $N|R|$ FLOPs in total.Consequently, the total FLOPs for the previous method is about $N(|R|+2|E|)+0.5|T|$.

Next we analyze the computation cost of \doph. For readers' convenience, we repeat the propagation scheme here:

During the computation, for each node $v^k$, we maintain a tuple $(v^k,{\mathbf{g}}^k,s^k):=(v^k, L\nabla v^k, \mathcal{L}v^k)$. The propagation rule of this tuple is:
\begin{align}
    \label{forward:func_elp1}
        v^j &= F_j(\{v^i:i\to j\}) \\
        {\mathbf{g}}^j &= \sum\limits_{i:i\to j}\pab{F_{j}}{v^i}{\mathbf{g}}^i \label{forward:func_elp2}\\
        s^{j} &= \sum\limits_{\substack{i,l\\i\to j\ l\to j}}\ppabc{F_{j}}{v^{i}}{v^{l}}{\mathbf{g}^{i}}^{\top}D \mathbf{g}^{l} + \sum\limits_{i:i\to j} \pab{F_{j}}{v^{i}}s^{i}\label{forward:func_elp3}
\end{align}
    
For the proposed \doph method, we perform forward propagation along $\mathcal{G}$ to obtain $\phi(\rvx)$, $L\nabla\phi(\rvx)$ (recall that $L$ is a matrix that comes from the decomposition of $A$) and $\mathcal{L}\phi(\rvx)$. The second term $L\nabla\phi(\rvx)$ is calculated along $\mathcal{G}$ according to \cref{forward:func_elp2}. The third term, i.e., the target differential operator, is calculated along $\mathcal{G}$ according to \cref{forward:func_elp3}.

The computational cost of the \doph method is dominated by \cref{forward:func_elp2} and \cref{forward:func_elp3}. As previously discussed, \cref{forward:func_elp2} takes $N|E|$ FLOPs.
For the first term in \cref{forward:func_elp3}, we decompose its calculation into two steps. First, we compute $\{\rvg^{i\top} D \rvg^l\}_{i\leq l,\ (i, l)\in R}$. Since $D$ is a diagonal matrix only with diagonal element in $\{0,\pm1\}$, this computation takes $0.5\mathrm{rank}(D)|R|\leq 0.5 N|R|$ FLOPs in total.
Next, following the topological order of $\mathcal{G}$, we sum over $i\leq l$ for each $j$, deriving the first term in \cref{forward:func_elp3}. By leveraging the symmetry of the Hessian matrix and Gram matrix, we reduce this computation by a factor of 2, which is $0.5|T|$ FLOPs. The computational cost of the second term is negligible compared with the first term since $s^{i}$ is a scalar. 

Summing the computational cost of all the terms, we have that the \doph method uses at most $0.5N(|R|+2|E|)+0.5|T|$ FLOPs. 

In practice, a large percentage of operations are linear transformations, and for any linear operation $F_j$, $\ppabc{F_j}{v^i}{v^l}=0$ for any $i\to j,\ l\to j$. This means the value $|T|$ is much smaller than $N|R|$ and $N|E|$. Thus, our method is about two times faster than the previous \hes for computing second order differential operators of general neural network functions. 

\clearpage
\section{Case Study for MLP}\label{apdx_b_1}
In this section, we follow the notation in \cref{apdxA} for an MLP $\phi$ and show that there is further speedup comparing with the general 2x result stated in \cref{thm1}. 

For MLP, We could explicitly compute 
\bgeq{
|E|=\sum_{l=0}^L N_lN_{l+1},\ |T|\leq\sum_{l=0}^L N_{l+1}N_l(N_l-1),\ |R|=\sum_{l=0}^LN_l(N_l-1).
}

To compute the first term in \cref{forward:func_elp3}, note that
\begin{align}
   \pab{F_{k,i}}{u^k_j} &=\sigma'(W^k\rvu^k +b^k)_iW^k_{ij}\\
   \ppabc{F_{k,i}}{u^k_j}{u^k_l} &=\sigma''(W^k\rvu^k +b^k)_iW^k_{ij}W^k_{i,l},
\end{align}
so we actually have
\begin{align}
    &\sum_{j,l}\ppabc{F_{k,i}}{u^{k}_{j}}{u^{k}_{l}}(L \nabla u^{k}_{j})^\top D (L\nabla u^{k}_{l})\\
   =&(\frac{\sigma''}{\sigma'^2}(W^k\rvu^k +b^k))_i (L\nabla 
   u^{k+1}_i)^\top D  (L\nabla u^{k+1}_i).
\end{align}
This gives an alternative way to compute the first term in (\ref{forward:func_elp3}), whose total computation cost is reduced from $\mathrm{r}(D)|R| $ to $\mathrm{r}(D)\sum_{l=0}^LN_{l+1}$ FLOPs, where $\mathrm{r}$ stands for $rank$.

Recall that the first term of (\ref{forward:func_elp3}) is one of the dominant calculations, this manner certainly boost the efficiency.

\clearpage
\section{Proof of \cref{thm2}}\label{apdx_c}

We focus on the peak memory usage caused by the forward-mode Jacobian calculation, i.e., the storage of $\nabla v^i$(or $\rvg^i$), which usually dominates the memory cost. In a typical forward-mode Jacobian calculation, we write $\nabla v^i$ into memory when the algorithm is applied to $v_i$ and remove it from memory when all the direct subnodes of $v^i$ have been computed.

We denote
\bgeq{
\tau(i):=\max\{j:i\to j\}.
}
Then at the moment our \doph forward propagation comes to the node $v^j$, the memory consumption is
\bgeq{\label{eq:25}C(j):=N\sum_{i:i\leq j\leq \tau(i)}1.}
If we further denote $\mathcal{M}_1$ to be the peak memory usage when executing the forward-mode \doph, we have 
\bgeq{
\mathcal{M}_1=\max\limits_{j}C(j).
}
From \cref{eq:25} we clearly see that $\mathcal{M}_1\leq N|V|$, here $V$ is the set of nodes in the computation graph $\mathcal{G}$. Furthermore, for any $j<M,\ C(j)\leq N( |V|-1)$. As a result, $\mathcal{M}_1=N|V|$ if and only if every nodes is pointing to the end node $v^M$, which corresponds to either one-layer linear model or a multivariate elementary function taking in input $\rvx$ and directly giving the output.
These extreme neural network functions rarely occur in deep learning literature.

By contrast, in the Hessian calculation in AutoDiff, the computation graph becomes a combination of $\mathcal{G}$ and $\hat{\mathcal{G}}$. As suggested by \cref{eq: hessian second}, the node $\hat{v}^i$ is a direct subnode of $v^i$. Consequently, at the moment when $\nabla v^M$ is written into memory, every $\nabla v^i$ for $v^i \in \mathcal{G}$ have been written into memory and could not be released since $\hat{v}^i$ have not been computed yet. As a result, the peak memory usage(denoted as $\mathcal{M}_2$) is strictly larger than $N|V|$, and therefore larger than the peak memory of Forward Laplacian.

In the specific case when $\phi$ is MLP. For any node $u_i^k,\ k\in\frac 1 2\mathbb{N}$, 
\bgeq{\tau(\#u_i^k)=
\begin{cases}
\#u^{k+0.5}_{N_{k+1}},\ k=l \\
\#u^{k+0.5}_{i},\ k=l+0.5,\ \ l\in\mathbb{N},
\end{cases}
}
where $\#$ denotes the order index of the node representing certain neurons in $\mathcal{G}$.
Thus,
\bgeq{
\frac 1 N C(\#u_i^k)=
\begin{cases}
    N_{l}+1,\ k=l\\
    N_{l-1}+i,\ k=l+0.5,\ \ l\in\mathbb{N}.
\end{cases}
}
Thus we conclude that $\mathcal{M}_1\leq  N\max\limits_{l}N_l+N_{l-1}\lesssim N \frac 2 L \sum_l N_l=N\frac 2 L |V|\leq \frac 2 L \mathcal{M}_2.$
\qed
\clearpage
\section{Experiments Settings\label{apdx:exp_details}}
\paragraph{Hardware.} All the results are evaluated on a single NVIDIA Tesla V100 GPU.
\paragraph{Network structure.}
We use the MLP and MLP with Jacobian sparsity to benchmark different methods. The MLP with Jacobian sparsity means we split each data into some small blocks\\
$\rvx=(\rvx_1,\rvx_2,...\rvx_k)$ and independently operate each block with an MLP. The output is a sum of product of each MLP output, i.e.
\[\text{output}=\sum_{d}\prod_{i=1}^k[\text{MLP}^i(\mathbf{x}_i)]_d.\]
Here $i$ refers to the index of block and $d$ refers to the index of the output in each MLP. Detailed hyperparameters can be found in \cref{tab:MLP_structure}.
\begin{table}[h]
\caption{Hyperparameters\label{tab:MLP_structure}}
\begin{center}
\newcolumntype{Y}{>{\centering\arraybackslash}X}
\begin{tabularx}{\textwidth}{@{}lYY@{}}
\toprule\midrule
&MLP& MLP with Jacobian sparsity \\ \midrule
hidden dimension & 256 & 256 \\ 
input dimension & 64 & 64 \\
\#layer & 8 &   8  \\
\#blocks & - & 16 \\ 
output dimension for each MLP & - & 8 \\
\bottomrule
\end{tabularx}
\end{center}
\end{table}

\paragraph{Coefficient matrix used for each experiments.} For the MLP structure, the coefficient matrix $A$ is a $64\times64$ matrix. For the MLP with Jacbobian sparsity, the coefficient matrix $A$ is a $64\times 64$ block diagonal matrix. The coefficient matrices are listed in \cref{tab:coefficient}. Here, $(\alpha_{ij})_{i,j}$ is a $64\times 64$ matrix and $(\sigma_{ij})_{i,j}$ is a $4\times 4$ matrix. Both $\alpha_{ij}$ and $\sigma_{ij}$ are drawn from a standard normal distribution. $(\delta_{ij})_{i,j}$ is the identity matrix such that $\delta_{ij}=0\text{ if }i\neq j\text{ else } 1$. $s_i=-1\text{ if }i=0 \text{ else } 1$. .

\begin{table}[h]
\caption{Coefficient matrix\label{tab:coefficient}}
\begin{center}
\newcolumntype{Y}{>{\centering\arraybackslash}X}
\begin{tabularx}{\textwidth}{@{}lYYY@{}}
\toprule\midrule
 Structure& Elliptic & Low-rank & general\\ \midrule
 MLP& $a_{ij}=\sum_{k=1}^{64} \alpha_{ik}\alpha_{jk}$ & $a_{ij}=\sum_{k=1}^{32} \alpha_{ik}\alpha_{jk}$&$a_{ij}=\delta_{ij}s_i$\\
 \\
 MLP with sparsity& $a_{il,jm}=\delta_{lm}\sum_{k=1}^{4} \sigma_{ik}\sigma_{jk}$ & $a_{il,jm}=\delta_{lm}\sum_{k=1}^{2} \sigma_{ik}\sigma_{jk}$&$a_{il,jm}=\delta_{lm}\delta_{ij}s_i$\\
\bottomrule
\end{tabularx}
\end{center}
\end{table}

\end{document}